News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 11 April 2003 10.00am EST

During the last week Chandra completed the observing schedule as planned.

The IRU-1 gyro-1 current has continued to display elevated and noisy behavior above its yellow limit (150mA) but below its red limit (200mA). Following discussions with the vendor, procedures and products for a B-side IRU swap are being readied in the event that they are needed. The gyro bias level has remained unaffected by the change in current and so to date, there has been no impact to science or operations.

A real-time procedure was completed on Apr 7 that verified the CPE warm start count has remained unchanged at 1 since Oct 18.

The observing schedule for next week is shown below and includes 2 observations of Saturn and observations of XTE J1908+094, 1136-135, and Cyg X-1 coordinated with the VLA, HST and RXTE respectively.

------------------------------------------
J2044+7044             ACIS-S       Apr 13
FIRST J142114+282452   ACIS-S
M87                    ACIS-S       Apr 14
Saturn (2 obs)         ACIS-S
Radiation Belts                     Apr 15
F12112+0305            ACIS-S
1136-135               ACIS-S
RCS0439.6-2905         ACIS-S       Apr 16
Radiation Belts                     Apr 17
Abell 3526 (2 obs)     ACIS-S       Apr 18
A1736                  ACIS-I       Apr 19
SDSS J145123.02-625.9  ACIS-S
XTE J1908+094          ACIS-S
Cyg X-1                ACIS-S/HETG
SDSS J143352+022714    ACIS-S       Apr 20
A2063                  ACIS-I

------------------------------------------

All spacecraft subsystems, with the exception of the MUPS thruster A-1 under investigation and gyro-1 as noted above, continued to operate nominally.

Return to Status Reports